OpenStudy (anonymous):

Choose the correct simplification of the expression a to the 7th power times b to the 8th power all over a to the 4th power times b to the 4th power. a11b12 a3b4 1 over a to the 3rd power times b to the 4th power 1 over a to the 3rd power times b to the 2nd power

OpenStudy (studygurl14):

\(\Large \frac{a^7b^8}{a^4b^4}\) Rule of exponents: \(\Large \frac{x^n}{x^m}=x^{n-m}\)